OpenStudy (anonymous):

Write the equation of direct variation that includes the point (-6, 2). @shamil98

OpenStudy (shamil98):

formula is y = kx

OpenStudy (shamil98):

i'm guessing you want to know what k is right? k = y/x

OpenStudy (anonymous):

So it would be k = 2/-6 ?

OpenStudy (solomonzelman):

In other word: Find the equation of a line that is going through (0,0) and (-6,2)

OpenStudy (shamil98):

k = 2/-6 = -1/3 plug in back to the equation y =-1/3x

OpenStudy (anonymous):

I'm sorry I ask so many questions, I'm just way behind. But you been a BIG help, thank you xD

OpenStudy (anonymous):

@SolomonZelman Thank you, as well c:

OpenStudy (solomonzelman):

Oh, Anytime!

OpenStudy (shamil98):

you're welcome